OpenStudy (anonymous):

If dy/dx = 7x^2/y^3 and y(3)=2, find an equation for y in terms of x

OpenStudy (amistre64):

separate the variables might help: dy/dx = 7x^2/y^3 (S) y^3 dy = (S) 7x^2 dx (1/4) y^4 = (7/3)x^3 + C y^4 = (28/3)x^3 + C y = +- 4root[(28/3)x^3 + C] maybe we can use logs to simplify?? )

OpenStudy (amistre64):

e^(y^4) = e^[(28/3)x^3 + C] e^4y = e^[(28/3)x^3] * e^C ; e^C is a constant.

OpenStudy (amistre64):

ln both sides... 4y = C * (28/3)x^3 y = C * 28x^3/12 y = C * (7x^3)/3 is the best I get.... if its right :)

OpenStudy (anonymous):

Ok I got it thanks!

OpenStudy (amistre64):

you got the answer? if so, was I right?

OpenStudy (anonymous):

Ya it was right

OpenStudy (amistre64):

Yay!!! thanx :)